Given the total cost, TC = 4 +Q, and the demand function is P = 10 – 2Q:

(a) Obtain the total revenue function.

(b) Find the profit function in terms of Q.

(c) Find the quantity when the profit is maximized. What is the maximum profit?


1
Expert's answer
2021-06-23T17:45:30-0400

(a)


"TR=pQ=(10-2Q)Q"

"TR=10Q-2Q^2"



(b)


"P=TR-TC=10Q-2Q^2-(4+Q)"

"P=-2Q^2+9Q-4"

(c)


"P'(Q)=-4Q+9"

Critical number(s)

"P'(Q)=0=>-4Q+9=0"

"Q=2.25"

"P(2)=-2(2)^2+9(2)-4=6"

"P(3)=-2(3)^2+9(3)-4=5"

The Profit has the absolute maximum with value of "6" when the quantity "Q=2."